# ERX: A Fast Real-Time Anomaly Detection Algorithm for Hyperspectral Line-Scanning

### v1

@@ IV-C2 Speed Comparison
This experiment compares the processing speed and scalability of the detection algorithms across increasing bands. Each algorithm is tested on the same randomly generated dataset of 500 pixels, 3,000 lines and 10 to 200 bands. The metric is the number of lines processed per second, averaged over 5 iterations. The results are shown in Figure 9.
Figure 9: A comparison all of the algorithms’ speed on the Jetson device with an increasing number of bands.
IV-C3 Detection Comparison
This experiment compares the anomaly detection performance of each algorithm. Each algorithm is tested on each of the three datasets. The AUC and F1 score are used again to evaluate detection performance. The average processing speed is recorded across 5 iterations. The results are shown in Table III.
TABLE III: Algorithm Detection Results for All Datasets (lps = lines per second).
Beach Dataset Synthetic Dataset SNP Dataset
Algorithm F1 score AUC Speed (lps) F1 score AUC Speed (lps) F1 score AUC Speed (lps)
RX Baseline 0.275 0.778 15 0.074 0.531 8 0.282 0.988 48
RT-CK-RXD 0.419 0.805 <1 0.067 0.581 14 0.391 0.992 19
RX-BIL 0.419 0.803 73 0.070 0.566 64 0.377 0.992 32
LBL-AD 0.192 0.697 14 0.047 0.585 386 0.576 0.971 658
ERX 0.441 0.845 273 0.150 0.852 248 0.306 0.986 886
V Discussion
V-A Results Discussion
Varying the momentum of ERX demonstrates that it can be tuned to improve detection performance in different scenarios (Table II). On the beach dataset, ERX performs the best for both very low momentum values (close to 0, which maintains the initial mean and covariance estimates from the first 99 lines which have no anomalies) and relatively high ones (closer to 1 which mostly uses the mean and covariance of the most recent line). Using momentum values between 1e-4 and 0.1 gives poorer detection performance for this dataset, because ERX is weighing anomalous pixels too heavily. Given that the beach dataset has a higher proportion of anomalous pixels, the heavier weighting leads to poorer background mean and covariance estimates.
The binary detection masks in Figure 8 indicate that smaller momentum values hinder ERX’s detection of the smaller targets at the end of the synthetic dataset, suggesting insufficient adaptation to the changing background. ERX performs best on the synthetic dataset for high momentum values, confirming that faster updates are advantageous when the scene distribution changes. This aligns with the results of the more homogeneous SNP dataset, where the best performance is for small momentum values. The consistently high AUC of the SNP dataset is expected from the high class imbalance.
However, $\alpha=0.01$ causes ERX to miss larger targets (larger targets contribute too much to the background), while $\alpha=0.001$ and lower miss smaller targets (the exponential decay of previous targets is too slow). Thus, momentum optimisation also depends on the frequency and size of anomalies, and not just on the variability of the scenery.
@@ V-A Results Discussion
The speed comparison (Figure 9) shows that ERX and LBL-AD are consistently faster than RX-Baseline, RX-BIL and RT-CK-RXD. Although LBL-AD is initially the fastest, ERX maintains the best performance as the number of bands moves past 100. The sharp drop in LBL-AD’s performance was found to occur upstream of Python, where the OpenBLAS library [40] for linear algebra computations switches from single-threading to multi-threading after a dimension size of 95 (i.e., greater than 95 bands). This issue arises when calling the Numpy dot product, which ERX avoids to maintain faster speeds at higher dimensions compared to the other algorithms.
Putting the results in the context of real-time operations, a modified OpenHSI camera that collected the beach dataset (108 band) operated at 120 lines per second with the Jetson. ERX, processing at 273 lines per second, is the only algorithm fast enough for real-time anomaly detection on the Jetson. It is also the fastest algorithm for the Sequoia National Park dataset.
The comparison in Table III shows that ERX outperforms in all metrics in the beach data set. RX-BIL and RT-CK-RXD detect comparatively well but are too slow for real-time processing. RT-CK-RXD failed to complete on the Jetson due to its pixel-wise approach and the OpenBLAS threading constraint. ERX is over three times faster than the next best algorithm, RX-BIL, on the Jetson.
ERX achieves the best detection scores and is the second fastest for the synthetic dataset. However, the performance of all algorithms is poor. The low F1 scores indicate many false detections and missed anomalies, and AUC scores near 0.5 mean that most algorithms are only slightly better than a random guessing. This shows that the synthetic data set is challenging. The high fraction of anomalous pixels compared to the beach and SNP datasets complicates separation from the background. The detection maps in figure 8 show that despite the low F1 score of ERX, it successfully identifies the structure of most targets. Therefore, object-based detection methods and metrics may be more practical than pixel-based approaches.
ERX has the fastest speed for the Sequoia National Park dataset but underperforms in detection. Reducing the momentum to $41e-$ 4 improves the F1 score, indicating $0.5$ updates ERX too quickly. Given the results of the speed test (Figure 9), one would expect LBL-AD to be faster than ERX at only 13 bands. This is a result of LBL-AD having a larger computational overhead on intialisation. However, LBL-AD excels in anomaly detection for this dataset, suggesting that it’s PCA approach works well with fewer bands. Further work is needed to confirm this hypothesis, and integrating similar dimensionality reduction techniques into ERX might enhance results.
V-B Limitations & Future Work

### v2

@@ IV-D1 Speed Test
This experiment compares the processing speed and scalability of the detection algorithms. Each algorithm is first tested on the same randomly generated dataset of 500 pixels, 3,000 lines, and an increasing number of bands between 10 to 200. The algorithms are then tested on another generated dataset with 3,000 lines, 50 bands, and an increasing number of pixels between 100 and 1,500. Both tests are repeated five times to get the average number of lines processed per second (LPS).
IV-D2 Detection Test
This experiment assesses the anomaly detection performance of each algorithm on the three datasets. Each model is also tested on the flipped version of each dataset (i.e. processed from right to left instead of left to right) to further assess the adaptability of each algorithm. This is particularly relevant for the beach and synthetic datasets, which start and end with different terrains. The tests are repeated five times to obtain average results and standard deviations, indicating model variability.
IV-D3 Ablation Study
This study analyses how altering or removing sparse random projection and the exponentially moving mean and covariance affect ERX’s performance. The first part adjusts the number of dimensions ($d$) reduced by sparse random projection: from 1 to 50 for the beach and synthetic datasets, and from 1 to 10 for the SNP dataset (since it only has 13 bands). Sparse random projection is also completely removed for all datasets. ROC curves will be generated for all datasets, and the speed will also be assessed because of its dependence on the number of dimensions. Only AUCs and speeds are provided for the synthetic dataset, as it presents the greatest detection challenge and changing $d$ has a consistent impact on speed regardless of the dataset.
The second part of the ablation study investigates the ERX momentum parameter $\alpha$, which modifies ERX’s detection performance based on the size of the anomaly. Different values are tested between $\alpha=1$ (using only the mean and covariance of the current line for detection) and $\alpha=1e-4$ (heavily weighing past data). This component is also eliminated by using incremental mean and covariance (IMC) updates. The implementation of the line-by-line IMC follows a batch version of Welford’s algorithm for online mean and covariance updates [81], which weighs all previous lines equally.
V Results & Discussion
V-A Speed Comparison
Figure 9 displays the results of the speed comparison: the top graph illustrates the speed changes with increasing bands, while the bottom graph shows the speed changes with increasing pixels per line (i.e., increasing line width).
Figure 9: A comparison of the algorithms’ average speed on the Jetson device across five iterations. The top graph shows speed versus the number of bands, and the bottom graph shows speed versus the number of pixels per line. The y-axis (speed) is on a log-10 scale, and the shaded areas represent the standard deviation.
@@ V-A Speed Comparison
Initially, ERX and LBL-AD are consistently faster than RX-Baseline, RX-BIL and RT-CK-RXD as the number of bands increases. However, ERX maintains the best performance as the number of bands moves past 95. The sharp drop in LBL-AD’s performance was found to occur upstream of Python, where the OpenBLAS library for linear algebra [82] switches from single-threading to multi-threading after a dimension size of 95. This issue arises when calling the Numpy dot product, which ERX bypasses to retain higher speeds than the other algorithms. RX-BIL and RX Baseline experience slight performance changes, whereas RT-CK-RXD becomes impractical beyond 95 bands. Putting the results in the context of real-time operations, a variant of the camera that collected the beach dataset (108 band) operated at 120 lines per second with the Jetson. ERX, processing at 561 lines per second (Table I), is the only algorithm fast enough for real-time anomaly detection. It is also the fastest algorithm for the synthetic dataset and the second fastest for the SNP dataset.
(a) Beach dataset ground truth
(b) RX Baseline
(c) RT-CK-RXD
(d) RX-BIL
(e) LBL-AD
(f) ERX
(a) Synthetic dataset ground truth
(b) RX Baseline
(c) RT-CK-RXD
(d) RX-BIL
(e) LBL-AD
(f) ERX
(a) SNP dataset ground truth
(b) RX Baseline
(c) RT-CK-RXD
(d) RX-BIL
(e) LBL-AD
(f) ERX
Figure 10: Anomaly heatmaps from each model for beach (top), synthetic (middle), and SNP (bottom) datasets. The top 10 pixels of each line are excluded from all beach heatmaps due to large radiometric distortions skewing the colour bar.
Figure 9 shows the RX-Baseline, RX-BIL, and RT-CK-RXD speeds rapidly decrease with increasing line width, while ERX and LBL-AD remain fast. There is a minor initial variation for LBL-AD attributed to processing noise on the Jetson, given that performance remains generally consistent for the remainder of the graph. In general, ERX is demonstrated as the most stable and scalable algorithm in this experiment.
V-B Detection Comparison
Figure 10 displays the models’ anomaly heatmaps for the original datasets. Figure 11 shows the mean ROC curves of all models for the original datasets and their flipped versions. The aggregated performance metrics are presented in Table I.
(g) Beach
(h) Synthetic
(i) SNP
(j) Beach (Flipped)
(k) Synthetic (Flipped)
(l) SNP (Flipped)
Figure 11: Mean ROC curves for all models across each dataset (both normal and flipped). The shaded area shows the standard deviation across five iterations.
TABLE I: Detection Results Across Five Iterations
Beach Speed Beach (Flipped)
Model AUC $\pm$ SD $\text{AUC}_{\text{TD}}$$\pm$ SD $\text{AUC}_{\text{BS}}$$\pm$ SD LPS AUC $\pm$ SD $\text{AUC}_{\text{TD}}$$\pm$ SD $\text{AUC}_{\text{BS}}$$\pm$ SD
RX Baseline $0.778\pm 0.000$ $0.821\pm 0.000$ $0.482\pm 0.000$ $12$ $0.774\pm 0.000$ $0.820\pm 0.000$ $0.480\pm 0.000$
RT-CK-RXD $0.805\pm 0.000$ $0.826\pm 0.000$ $0.515\pm 0.000$ $<1$ $0.795\pm 0.000$ $0.833\pm 0.000$ $0.498\pm 0.000$
RX-BIL $0.803\pm 0.000$ $0.828\pm 0.000$ $0.510\pm 0.001$ $62$ $0.792\pm 0.000$ $0.833\pm 0.000$ $0.494\pm 0.001$
LBL-AD $0.695\pm 0.000$ $0.792\pm 0.000$ $0.416\pm 0.000$ $12$ $0.710\pm 0.000$ $0.798\pm 0.000$ $0.428\pm 0.000$
ERX $\bm{0.941\pm 0.005}$ $\bm{0.890\pm 0.006}$ $\bm{0.659\pm 0.017}$ $\bm{561}$ $\bm{0.938\pm 0.002}$ $\bm{0.896\pm 0.002}$ $\bm{0.635\pm 0.008}$
Synthetic Speed Synthetic (Flipped)
Model AUC $\pm$ SD $\text{AUC}_{\text{TD}}$$\pm$ SD $\text{AUC}_{\text{BS}}$$\pm$ SD LPS AUC $\pm$ SD $\text{AUC}_{\text{TD}}$$\pm$ SD $\text{AUC}_{\text{BS}}$$\pm$ SD
RX Baseline $0.531\pm 0.000$ $0.699\pm 0.000$ $0.336\pm 0.000$ $4$ $0.527\pm 0.000$ $0.696\pm 0.000$ $0.334\pm 0.000$
RT-CK-RXD $0.581\pm 0.000$ $0.708\pm 0.000$ $0.379\pm 0.000$ $10$ $0.632\pm 0.000$ $0.695\pm 0.000$ $0.457\pm 0.000$
RX-BIL $0.566\pm 0.002$ $0.714\pm 0.001$ $0.357\pm 0.001$ $64$ $0.615\pm 0.001$ $0.724\pm 0.002$ $0.400\pm 0.002$
LBL-AD $0.576\pm 0.000$ $0.509\pm 0.001$ $\bm{0.563\pm 0.000}$ $361$ $0.550\pm 0.000$ $0.611\pm 0.000$ $0.433\pm 0.000$
ERX $\bm{0.751\pm 0.032}$ $\bm{0.807\pm 0.019}$ $0.463\pm 0.016$ $\bm{508}$ $\bm{0.760\pm 0.022}$ $\bm{0.764\pm 0.023}$ $\bm{0.514\pm 0.004}$
SNP Speed SNP (Flipped)
Model AUC $\pm$ SD $\text{AUC}_{\text{TD}}$$\pm$ SD $\text{AUC}_{\text{BS}}$$\pm$ SD LPS AUC $\pm$ SD $\text{AUC}_{\text{TD}}$$\pm$ SD $\text{AUC}_{\text{BS}}$$\pm$ SD
RX Baseline $0.988\pm 0.000$ $\bm{0.871\pm 0.000}$ $\bm{0.880\pm 0.000}$ $38$ $0.984\pm 0.000$ $0.870\pm 0.000$ $\bm{0.860\pm 0.000}$
RT-CK-RXD $\bm{0.992\pm 0.000}$ $\bm{0.871\pm 0.000}$ $0.864\pm 0.000$ $15$ $\bm{0.986\pm 0.000}$ $0.883\pm 0.000$ $0.815\pm 0.000$
RX-BIL $\bm{0.992\pm 0.000}$ $0.869\pm 0.000$ $0.857\pm 0.001$ $35$ $\bm{0.986\pm 0.000}$ $\bm{0.887\pm 0.001}$ $0.803\pm 0.001$
LBL-AD $0.971\pm 0.000$ $0.818\pm 0.003$ $0.814\pm 0.005$ $\bm{787}$ $0.968\pm 0.000$ $0.835\pm 0.002$ $0.779\pm 0.004$
ERX $0.959\pm 0.024$ $0.866\pm 0.006$ $0.754\pm 0.055$ $561$ $\bm{0.986\pm 0.002}$ $0.871\pm 0.005$ $0.809\pm 0.009$
Figure 11 shows that ERX is the overall top performer, excelling on the beach and synthetic datasets and remaining competitive on the SNP dataset, demonstrating its adaptability and robustness. ERX’s ROC curves are much more desirable, vastly outperforming the other algorithms on the HSI datasets even with model variation. Figure 10 shows that ERX uniquely highlights the black tarp in the beach dataset. Some brighter patches of vegetation and water areas in ERX’s heatmap implies higher sensitivity to local outliers. This reflects an adaptive trade-off, which loses some prior information due to the exponential moving mean and covariance. RX-BIL and RT-CK-RXD effectively highlight most targets, but LBL-AD misses most, detecting only a tarp and cars, indicating a loss of essential spectral information in the real-time PCA technique.
Table I further highlights the strong performance of ERX on the beach dataset, excelling across all detection metrics and in speed. ERX achieves high mean AUCs for both normal and flipped images, demonstrating its adaptability to changing scenery and different initial backgrounds. The larger margin between ERX and other algorithms’ $\text{AUC}_{\text{BS}}$ score highlights its primary strength in reducing false alarms, suggesting its adaptive methodology excels at background suppression. RX-BIL and RT-CK-RXD detect well but are too slow for real-time use. RT-CK-RXD is especially slow on the Jetson due to the OpenBLAS threading limitations and its pixel-wise approach. ERX is over nine times faster than RX-BIL, the next-fastest algorithm. ERX’s scalability, adaptability, and robustness are demonstrated given the high-dimensionality, the changing scenery, and the geometric distortions.
(a) Beach
(b) Synthetic
(c) SNP
Figure 12: Average ROC curves for ERX with different dimensions using sparse random projection. Shaded areas indicate the standard deviation over ten iterations.
ERX also achieves the best results on the synthetic dataset. ERX’s adaptability is evident with its high mean AUC between normal and flipped images, unlike RX-BIL and RT-CK-RXD which show significant changes. In the flipped dataset the smaller anomalies are processed first, contributing less to initial background estimates and enhancing background-anomaly separation for these two algorithms. All algorithms struggle to perform well on the this dataset given that most have AUC scores near 0.5, which is only slightly better than random guessing. It is challenging due to a higher proportion of anomalous pixels, complicating anomaly-background separation compared to the beach and SNP datasets. The higher standard deviations for ERX suggest that the five dimensions ($d=5$) used in sparse random projection do not fully capture all the information needed to consistently separate anomalies.
The high frequency of anomalies affects the RX Baseline and LBL-AD algorithms the most, as their ROC curves fall below the random model at lower thresholds (top right of Figures 11 and 11). The impact is visualised in the anomaly heatmaps, where LBL-AD mistakenly detects the background as anomalous. Overall, ERX performs well and Figure 10 shows that it detects most panel structures despite a generally low mean AUC. ERX’s consistently higher $\text{AUC}_{\text{TD}}$ scores support this, but the relatively low $\text{AUC}_{\text{BS}}$ scores for all algorithms highlight the general challenge of background suppression in this dataset. Object-based detection methods that factor in the spatial structure of neighbouring anomalies could be more practical than pixel-based approaches.
ERX has the second fastest speed for the Sequoia National Park dataset but slightly underperforms the other algorithms for mean AUC (Table I). LBL-AD is the fastest on this lower-dimensional dataset, which is consistent with the findings in the speed comparison. All algorithms perform well on this dataset, which can be attributed to the stable background with minimal change in scenery. Visually, LBL-AD effectively highlights anomalies but exhibits a brighter background (Figure 10), indicating sensitivity to noise. This aligns with the hard assumption of a homogeneous background. RX Baseline, RT-CK-RXD, and RX-BIL separate anomalies well, while ERX generally performs well but struggles slightly with background suppression. This can be seen by the brighter spots in the anomaly heatmap (similar to the beach dataset) and the lower $\text{AUC}_{\text{BS}}$ score. Reducing the momentum for slower updates should reduce this noise.
ERX has a larger model variance for the SNP dataset compared to the flipped counterpart and the other datasets, shown in Figure 11 and by the AUC standard deviation (Table I). The reason is unclear, but the small number of iterations (five) could be the cause. More iterations could stabilise model performance. However, ERX still maintains strong absolute detection and a high speed, achieving AUCs above $0.959$ on both normal and flipped datasets.
V-C Ablation Study
V-C1 Sparse Random Projection
Figure 12 shows the ROC curves of ERX with varying projected dimensions for each dataset, and the detection performance and speed results of the synthetic dataset are shown in Table II.
TABLE II: Ablation Study - Varying Projected Dimensions ($d$) on the Synthetic Dataset.
$d$ AUC $\pm$ SD Speed (LPS)
$1$ $0.750\pm 0.085$ $\bm{610}$
$3$ $\bm{0.783\pm 0.014}$ $580$
$5$ $0.759\pm 0.027$ $566$
$10$ $0.698\pm 0.012$ $525$
$20$ $0.596\pm 0.004$ $446$
$30$ $0.532\pm 0.005$ $379$
$50$ $0.452\pm 0.002$ $287$
No SRP $0.356\pm 0.000$ $243$
Increasing the number of projected dimensions ($d$) reduces the variance of ERX as more spectral information is retained from the original image after SRP. For the beach and synthetic datasets, originally with 108 and 90 bands, most of the information is retained after 10 dimensions. Balancing the number of projected dimensions is crucial, as it also reduces the processing speed exponentially (Table II). ERX is already a fast algorithm due to the quick operation of exponentially moving averages, yet removing SRP altogether clearly shows that it substantially enhances the speed further (by up to 2.5 times on the synthetic dataset, as an example).
However, excessively increasing dimensionality also reduces the detection performance on these datasets. This is seen in the deterioration in the ROC curves and the mean AUC when $d>\approx 5$. This can be expected because of the high correlation between hyperspectral bands and the curse of dimensionality, where too many features hinder anomaly-background separation. On the synthetic dataset especially, ERX becomes worse than a random guessing model after more than 30 dimensions. This suggests that anomalies and background pixels share similar values across many bands, leading to significant redundancy. AUCs below 0.5 indicate that the algorithm is modelling anomalies as the background, which is exacerbated by the high frequency of the anomalies. Reducing dimensionality via SRP mitigates this issue.
TABLE III: Ablation Study - Mean AUC $\pm$ SD with Varying Momentum ($\alpha$).
$\alpha$ Beach Beach (Flipped) Synthetic Synthetic (Flipped) SNP SNP (Flipped)
$1$ $0.933\pm 0.006$ $0.931\pm 0.003$ $0.725\pm 0.020$ $0.700\pm 0.035$ $0.977\pm 0.021$ $0.952\pm 0.055$
$0.9$ $0.931\pm 0.004$ $0.931\pm 0.004$ $0.725\pm 0.024$ $0.704\pm 0.027$ $0.964\pm 0.054$ $0.974\pm 0.021$
$0.5$ $0.935\pm 0.004$ $0.933\pm 0.006$ $0.732\pm 0.034$ $0.700\pm 0.034$ $0.956\pm 0.061$ $0.964\pm 0.046$
$0.1$ $0.937\pm 0.007$ $0.941\pm 0.004$ $\bm{0.773\pm 0.013}$ $0.777\pm 0.029$ $0.989\pm 0.001$ $\bm{0.980\pm 0.017}$
$0.01$ $\bm{0.956\pm 0.005}$ $\bm{0.956\pm 0.005}$ $0.735\pm 0.023$ $\bm{0.788\pm 0.016}$ $0.979\pm 0.022$ $0.925\pm 0.099$
$31e-$ 3 $0.845\pm 0.045$ $0.845\pm 0.034$ $0.551\pm 0.013$ $0.267\pm 0.034$ $0.959\pm 0.093$ $0.959\pm 0.078$
$41e-$ 4 $0.779\pm 0.027$ $0.746\pm 0.015$ $0.552\pm 0.017$ $0.330\pm 0.071$ $0.982\pm 0.015$ $0.974\pm 0.024$
No EMAs $0.922\pm 0.025$ $0.845\pm 0.033$ $0.720\pm 0.020$ $0.623\pm 0.102$ $\bm{0.990\pm 0.003}$ $0.978\pm 0.026$
Unlike the hyperspectral datasets, the SNP dataset shows much larger standard deviations at low band numbers, even though it originally contains only 13 bands. Detection performance consistently improves and stabilises with increased dimensions instead of degrading after 5 or 10. This is a result of the sparsity condition within SRP, whereby only $\frac{1}{\sqrt{b}}$ of the weights in the randomly generated projection matrix are actually active. Given three projected dimensions ($d=3$), ERX has around 28 and 31 active weights for synthetic and beach datasets respectively, while the SNP dataset only has around 18. Thus, using the recommended density of $\frac{1}{\sqrt{b}}$ for low-dimensional inputs risks losing core information as some key bands may not contribute at all. For lower-dimensional, less correlated data, removing SRP may be more beneficial if speed advantages do not outweigh the loss of information. Nonetheless, for high-dimensional hyperspectral data, sparse random projection to less than 10 dimensions has retained most of the spectral band information, while substantially improving the speed and detection performance of ERX.
V-C2 Momentum
The detection results for varying momentum across all datasets are given in Table III, and Table IV compares the average speeds when using exponential moving averages (EMAs) and when they are removed. Anomaly heatmaps of the synthetic dataset in Figure 13 visually demonstrate how momentum affects ERX’s detection performance.
TABLE IV: Ablation Study - Speed (Lines per Second) With and Without Momentum ($\alpha$).
Model Beach Synthetic SNP
No EMAs $527$ $535$ $574$
EMAs $\bm{573}$ $\bm{548}$ $\bm{599}$
(a) $\alpha=1$
(b) $\alpha=0.1$
(c) $\alpha=1e-3$
(d) No EMAs
Figure 13: ERX anomaly heatmaps of the synthetic dataset with varying momentum ($\alpha$). “No EMAs” replaces ERX’s exponentially moving averages with incremental mean and covariance updates, equally weighing all previous lines.
Adjusting the momentum of ERX demonstrates that it improves the detection performance in various scenarios (Table III). ERX works well with high momentum on the beach and synthetic datasets, but performance declines after $\alpha<0.01$. This suggests that when $\alpha\leq 1e-3$, ERX is slow to adapt and over-emphasises past pixels. Small $\alpha$ values overly bias the model toward initial background estimates given by the mean and covariance of the first line, further slowing the adaptation to changing scenery. The synthetic dataset and the flipped version are most affected, as the high anomaly rate further skews the background mean and covariance estimates. Figure 13 illustrates this, with anomalies blending into the background after the first two columns of anomalous targets. ERX performs consistently across different momentum values on the SNP dataset, as expected due to more uniform scenery and infrequent anomalies. Momentum values of 0.1 and 0.01 generally perform best across all datasets, but it is clear that adjusting the momentum can enhance anomaly detection based on the size and frequency of the anomalies in the scene.
Removing the exponentially moving averages and switching to the incremental mean and covariance (IMC) shows comparable performance in some instances, mostly in the less diverse SNP dataset. Figure 13 shows that the IMC still suppresses the background well but misses the third column of anomalies. This anomaly column is located where the background of transitions from vegetation to sand, highlighting an inability to adapt without EMAs. However, a notable difference from removing the EMAs is the reduced mean AUC when the beach and synthetic datasets are flipped. The equal-weighting approach of IMC clearly depends on the starting position, while EMA is more flexible and delivers consistent results. Overall, the exponentially moving averages are a vital part of ERX, and are crucial for adapting to changing scenery and outperforming equal-weight methods in speed and detection.
V-D Limitations & Future Work